Simplify the expression. Write the reslut using positive exponents only.
5^-7x^-3y^2
____________
5^-4x^-6y^4

Answer by checkley71(8403)

5^-7/5^-4=5^(-7+4)=5^-3=1/5^3
X^-3/X^-6=X^(-3+6)=X^3
Y^2/Y^4=Y^(2-4)=Y^-2=1/Y^2
COMBINING THESE ELEMENTS WE HAVE:
X^3/5^3Y^2 OR X^3/125Y^2 ANSWER.